OpenStudy (anonymous):

Simple question, why does 3--2=5???

OpenStudy (anonymous):

+

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

3+2 = 5

OpenStudy (anonymous):

Because of Piano's Axiom!

OpenStudy (anonymous):

This should help: 3 - (-2) = 3 -1*(-2) -1*(-2) = 2, so 3-(-2) = 3+2 = 5

OpenStudy (anonymous):

3 + 2 means the second succession of 3 which is 5.

OpenStudy (anonymous):

Because of the rule when two negative sign are next to each other it becomes positive/plus sign: -- = + -+ = - ++ = +

OpenStudy (anonymous):

Cheers guys. @ opiesche thumbs up!!

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

yes because two negatives make a positive